Space Stations

NASA International Space Station On-Orbit Status 3 January 2017

By Marc Boucher
Status Report
January 4, 2017
Filed under , ,
NASA International Space Station On-Orbit Status 3 January 2017
A satellite is ejected from the Japan Aerospace Exploration Agency (JAXA) Small Satellite Orbital Deployer on the International Space Station on Dec. 19, 2016. The satellite is actually two small satellites that, once at a safe distance from the station, separated from each other, but were still connected by a 100-meter-long Kevlar tether. Credit: NASA.

In a remarkable demonstration of robotic prowess, ground controllers used the Canadian-built “Dextre” Special Purpose Dexterous Manipulator over the weekend to install three new lithium-ion batteries in the International Space Station’s 3A power channel Integrated Electronics Assembly (IEA) pallet on the starboard 4 truss.
Dextre also removed four old nickel-hydrogen batteries from the IEA, three of which were stowed on the Japanese H-II Transfer Vehicle’s external pallet to wrap up the first act of a complex procedure to upgrade the station’s power system. A fourth old battery was temporarily stowed on a platform on Dextre.

This clears the way for the first of two spacewalks Friday in which Expedition 50 Commander Shane Kimbrough and Flight Engineer Peggy Whitson of NASA will install three adapter plates in slots on the IEA to which three of the old nickel-hydrogen batteries will be mounted to remain on the ISS but will be dormant. In all, nine nickel-hydrogen batteries will be stowed on the external pallet for disposal when the HTV is deorbited to burn up in the Earth’s atmosphere late this month.

Three additional new lithium-ion batteries flown to the ISS aboard the HTV will be robotically installed in the starboard truss’ 1A power channel Integrated Electronics Assembly between Friday’s spacewalk and a second spacewalk scheduled Jan. 13 for Kimbrough and Flight Engineer Thomas Pesquet of the European Space Agency. Five additional nickel-hydrogen batteries will be removed robotically from the IEA prior to the second spacewalk.

A briefing to preview the two spacewalks and to review all of the robotics work will be broadcast on NASA Television on Wednesday, Jan. 4 at 2 p.m. Eastern time.

On-Orbit Status Report

Extravehicular Activity (EVA) Preparations: The USOS crew continued preparations for Friday’s EVA by auditing and configuring EVA tools, reviewing EVA Cuff Checklist procedures and participating in a conference with ground specialist. Additionally, they checked out the Rechargeable EVA Battery Assemblies (REBAs) on Extravehicular Mobility Unit (EMU) suits 3006 and 3008 to verify the glove heaters are functional and helmet cameras are receiving power.

Mobile Servicing System (MSS) Operations: Yesterday and overnight (GMT 002-003), Robotics Ground Controllers powered up the MSS and operated the Space Station Remote Manipulator System (SSRMS) and the Special Purpose Dexterous Manipulator (SPDM) Body and Arm2 to remove a Nickel-Hydrogen (NiH2) battery from slot 3 on the 3A Integrated Electronics Assembly (IEA) and temporarily stow it on SPDM Enhanced ORU Temporary Platform (EOTP) Side 3. They then removed a Lithium-Ion (Li-Ion) battery from Slot C on the H-II Transfer Vehicle 6 (HTV6) Exposed Pallet (EP) and installed it in Slot 3 of the 3A Integrated Electronics Assembly (IEA), this was the last Li-Ion battery that had to be installed in the 3A IEA. The team is currently performing operations to drive the slot 1 and slot 5 H1 bolts with the ROST to complete the 3A pre-EVA robotics ops.

Fluid Shifts – Baseline Imaging: The last of the three 49 Soyuz subjects performed their Baseline Imaging measurements with assistance from both onboard operators and ground remote guiders. Ultrasound imaging was taken of arterial and venous measures of the head and neck, cardiac, ophthalmic and portal vein, and tissue thickness of lower and upper body. Additional measurements using the Cerebral and Cochlear Fluid Pressure (CCFP), Distortion Product Otoacoustic Emissions (DPOAE), Optical Coherence Tomography (OCT), ESA’s Cardiolab (CDL) Holter Arterial Blood Pressure Device, and Intraocular Pressure (IOP) using the tonometer were taken as well. Fluid Shifts is a joint NASA- Russian experiment that investigates the causes for severe and lasting physical changes to astronaut’s eyes. Because the headward fluid shift is a hypothesized contributor to these changes, reversing this fluid shift with a lower body negative pressure device is investigated as a possible intervention. Results from this study may help to develop preventative measures against lasting changes in vision and eye damage.

European Space Agency (ESA) Education Payloads Operations (EPO): The crew reviewed the overview for the three EPO activities which make up the EXo-ISS – CERES, CrISStal and CatalISS. They also installed the CERES experiment in the Veggie facility, fastening the Seed Support then watering the seeds. The crew then prepared the CrISStal Salt Bag, dissolving the Seingnette salt then injecting the solution into the experiment bags. Finally, the crew initiated the CatalISS experiment by activating the CatalISS Gelatin Syringe by injecting degassed lemon-lime soda. CERES, is devoted to the study of plant growth; lentil, mustard, and radish seeds were selected having met the required criteria of rapid germination, vertical growth, and resistance to long-term storage. CrISStal, aims to demonstrate the role of gravity in growing crystals, where in microgravity growth is the same in all directions, giving a perfect regular shape of the crystal. The crystal grows in a liquid solution from saturated Rochelle salt. The student designed experiment, CatalISS, aims to prove the influence of gravity on the catalytic reactions. For this purpose, the experiment operates from catalysis of enzymes and gelatin for reaction with a duration of three days.

Today’s Planned Activities
All activities were completed unless otherwise noted.

RБМП Ф1 absorption cartridge regeneration initiate
Unstow of the EPO Pesquet Pouch for the Ceres, Crisstal, and Cataliss experiments.
Fluid Shifts Ultrasound 2 HRF Rack 1 Power On
Pre-treat tank [Е-К] and hose R&R in АСУ system, АСУ activation post R&R
Fluid Shifts OCT Baseline Setup
Reading of Big Picture Words for Ceres, Crisstal, Cataliss
Fluid Shifts CDL Holter Arterial BP Hardware Don – Subject
Fluid Shifts Ultrasound Baseline Scan
Vozdukh [АВК СОА] valves service activation
БД-2 adjustment for physical exercise
Crew Medical Officer (CMO) Proficiency Training
Fluid Shifts Ultrasound Baseline Scan – Operator
Photo and Video Recording of Life on Station
Regenerative Environmental Control and Life Support System (RGN) WSTA Fill
PROFILAKTIKA-2. Assistance in experiment setup
ИП-1 sensors positioning check
Nikon Camera Time Sync
Fluid Shifts CCFP Baseline Test
PROFILAKTIKA-2. Closeout ops (assistance)
Fluid Shifts OCT Baseline Exam
Central Post Laptop Monthly Maintenance Log-files Downlink
PELIKAN PL Photo Spectral System Battery Charge (init)
Fluid Shifts DPOAE Baseline Test – Subject
Fluid Shifts OCT Baseline Stow
Fluid Shifts Tonometry Baseline Setup
Fluid Shifts Tonometry Baseline Exam
Fluid Shifts CDL Holter Arterial BP Hardware Doff – Subject
Formaldehyde Monitoring Kit (FMK) Deployment Operations
Grab Sample Container (GSC) Sampling Operations
Veggie Hardware and display OBT
Fluid Shifts Tonometry Baseline Stow
Filling (separation) of ЕДВ (КОВ) for Elektron or ЕДВ-СВ
Setup of the EPO Pesquet Ceres experiment
Fluid Shifts CCFP Baseline Stow
COTS UHF Communication Unit (CUCU) Circuit Breakers Close
Regenerative Environmental Control and Life Support System (RGN) WSTA Fill
СОЖ maintenance
Progress 433 Stowage Ops (DC1) w/IMS update
COTS UHF Communication Unit (CUCU) Circuit Breaker Open
Extravehicular Activity (EVA) Tool Configuring
Setup of the EPO Pesquet Crisstal Experiment.
Public Affairs Office (PAO) High Definition (HD) Config LAB Setup
PAO Preparation
Public Affairs Office (PAO) Event in High Definition (HD) – Lab
Setup of the Cataliss experiment
Extravehicular Mobility Unit (EMU) Cuff Checklist Print
Extravehicular Mobility Unit (EMU) Resize
CONTENT. Experiment ops
Rechargeable EVA Battery Assembly (REBA) Powered Hardware Checkout
CCE Preparation
Extravehicular Mobility Unit (EMU) Relocate
Extravehicular Activity (EVA) Procedure Conference
DUBRAVA. Observation and photography
PELIKAN Observation and photography using Photo Spectral System
БМП Ф1 absorption cartridge regeneration; (end)

Completed Task List Items
Commercial Generic Bioprocessing Apparatus 1, Compact Flash Card [Completed Saturday]
Crew Provisions Relocate [Completed Saturday]
LAB PS-120 label fix [Completed Saturday]
Configuring the EVA GoPro Camera [Completed Saturday]
COL1O2 Cleanup [Completed Saturday]
Payload Hardware Relocate [Completed Saturday]
Tropical Cyclone Ops Overview Review [Completed Saturday]
On-Orbit Hearing Assessment (O-OHA) with EarQ Software Setup and Test [Completed Saturday]
JEM CTB CONSOLIDATION [Completed Saturday]
WiseNet Base Station Antenna Check [Completed Saturday]
Cell Biology Experiment Facility (CBEF) Ventilation Valve Pedestal Installation [Completed Saturday]
Portable Emergency Provisions (PEPS) Inspection [Completed Sunday]
Total Organic Carbon Analyzer (TOCA) Water Recovery System (WRS) Sample Analysis [Completed Sunday]
EVA Simplified Aid For EVA Rescue Practice [Completed Sunday]
Lighting Effects Visual Performance Tests Light Setting [Completed Sunday]
Plug-in Plan Audit – NOD1 [Completed Sunday]
Children’s Artwork Calendar Downlink Message [Completed Sunday]
Columbus HD video Recording on SSDs [Completed Sunday]
Total Organic Carbon Analyzer (TOCA) Sample Data Record [Completed Sunday]
Emergency Ammonia Measurement Kit Spare Battery Changeout [Completed Sunday]
Recording of AstroPi Message [Completed Sunday]
ESA PAO Recorded Message “Opening of the « Journees de linnovation en sante »” [Completed Sunday]
Tomatoshpere Video Record [Completed Monday]
NOD1PD2 Light Installation
NOD1SD2 Light Installation

Ground Activities
All activities were completed unless otherwise noted.
CUCU Checkout
Day 4 – 3A Battery R&R Robotic Operations

Three-Day Look Ahead:
Wednesday, 01/04: Spheres Setup and Run, EVA tool and Camera Config, Body Measures
Thursday, 01/05: EVA Procedure Review, EVA Conference, Airlock Equipment Lock Prep, EVA Tool Audit
Friday, 01/06: Airlock Depress, S4 Battery EMU EVA

QUICK ISS Status – Environmental Control Group:
Component – Status
Elektron – Off
Vozdukh – Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) – Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) – On
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Operate
Major Constituent Analyzer (MCA) Node 3 – Idle
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Process
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full Up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.